OpenStudy (lionhowlter):

What is the slope of the line that passes through the points (3,1) and (-2,5) a.) 5/4 b.) -4/5 c.) 4,5 d.) -5/4

OpenStudy (fedorable):

you have an amazing username :)

OpenStudy (lionhowlter):

thank you so much :) are you in the phandom?

OpenStudy (samanthagreer):

Hi welcome to openstudy this isant the correct section for you question but dont worry i will help you find where you need to be up at the top were it says openstudy next to it you something that says openstudy feedback hover over it and click find more subjects and then ou can find any subject you need and since you are new i sugest reading this http://openstudy.com/code-of-conduct and also this http://openstudy.com/terms-and-conditions yes and it wouldnt hurt to read this one two http://openstudy.com/privacy-policy once agian welcome to openstudy HAPPY STUDING

pooja195 (pooja195):

Hi, although you are in the wrong section i'll help you ^_^ So for this we use the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] Can you plug in the numbers and find the slope?

OpenStudy (lionhowlter):

3-(-2) 1-5 i think

pooja195 (pooja195):

(x1,y1) (x2,y1) \[\LARGE \frac{5-1}{-2-3}=slope\] Ok now solve

OpenStudy (lionhowlter):

4/-1?

pooja195 (pooja195):

4 is right what is -2-3?

OpenStudy (lionhowlter):

i'm still getting -1?

OpenStudy (lionhowlter):

oh wait it's -5

OpenStudy (lionhowlter):

4/-5

pooja195 (pooja195):

Yes :) Correct well done ^_^

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (lionhowlter):

Thank you so much!!

pooja195 (pooja195):

Yw ^.^